OpenStudy (anonymous):

If the scale factor between two similar solids is 2:5, then the ratio of their lateral areas is 4:10

OpenStudy (anonymous):

will award medals

OpenStudy (anonymous):

true or false

OpenStudy (anonymous):

@SolomonZelman lz help

OpenStudy (anonymous):

@jim_thompson5910 plz help me

OpenStudy (anonymous):

somebody help me

jimthompson5910 (jim_thompson5910):

This is false because the ratio of the areas is really 4:25 Notice how 4 = 2^2 and 25 = 5^2

OpenStudy (anonymous):

he ratio for the lengths of radii of similar cylinders is 6 : 4. What is the ratio of their volumes?

OpenStudy (anonymous):

what about this @jim_thompson5910

jimthompson5910 (jim_thompson5910):

if two figures have ratio a:b, then their volumes are in ratio a^3:b^3

jimthompson5910 (jim_thompson5910):

it might help to reduce 6:4 first

OpenStudy (anonymous):

so 27 and 8 ??

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910 plz help

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

In a certain cylinder the sum of the areas of the two bases is equal to the lateral area. How is the length of the altitude related to the radius? what about this

OpenStudy (anonymous):

@jim_thompson5910 plz help

OpenStudy (anonymous):

@jim_thompson5910 plz help me

jimthompson5910 (jim_thompson5910):

Each base has an area of pi*r^2

jimthompson5910 (jim_thompson5910):

The lateral area is 2*pi*r*h

jimthompson5910 (jim_thompson5910):

" the sum of the areas of the two bases is equal to the lateral area" So, pi*r^2 + pi*r^2 = 2*pi*r*h solve for h to get your answer

OpenStudy (anonymous):

can u do that for me plz cause i am not good at it

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok this will be the last one I'll help with

OpenStudy (anonymous):

ok sir

jimthompson5910 (jim_thompson5910):

pi*r^2 + pi*r^2 = 2*pi*r*h 2*pi*r^2 = 2*pi*r*h pi*r^2 = pi*r*h ... Divide both sides by 2 r^2 = r*h ... Divide both sides by pi r = h ... Divide both sides by r So that means h = r (the height and radius are the same, whatever they are)